Find the mean, variance, and standard deviation of the binomial distribution with the given values of n and p. n=70, p=0.6 

The mean, μ, is _______  (Round to the nearest tenth as needed.) 

The variance, σ2, is _______ (Round to the nearest tenth as needed.) 

The standard deviation, σ, is _______ (Round to the nearest tenth as needed.)
Answer #1

Given that
n= 70,
p = 0.6

We can calculate the mean, variance and standard deviation as follows
Mean= μ= np = 70× 0.6 = 42
Mean = μ = 42.0
Variance = σ^2= [ np (1-p) ]
= [ (70×0.6)(1-0.6) ]
= (42× 0.4)
Variance = σ^2 = 16.8
Standard deviation= σ=✓variance
=✓16.8
=4.0988
Standard deviation= σ = 4.1 ...(As required to nearest 10th )